Find using distributive property: 
4275 x 125 ??????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given 4275 × 125
= (4000 + 200 + 100–25) × 125
= (4000 × 125 + 200 × 125 + 100 × 125–25 × 125)
= 500000 + 25000 + 12500–3125
= 534375

Thanks